OpenStudy (anonymous):

Suppose oil spills from a ruptured tanker and spreads in a circular pattern. If the radius of the oil spill increases at a constant rate of 1.5 m/s. How fast is the area of the spill increasing when the radius is 19 m?

OpenStudy (bahrom7893):

A = pi*r^2 dA/dt = 2pi*r

OpenStudy (anonymous):

so 2(pi)(19)^2?

OpenStudy (bahrom7893):

No wait i forgot the dr/dt

OpenStudy (bahrom7893):

dA/dt = 2pi*r*(dr/dt) = 2pi*19(m)*1.5(m/s)

OpenStudy (bahrom7893):

dA/dt = 57pi (m^2/s)

OpenStudy (bahrom7893):

As you can see, the units work out nicely :)

OpenStudy (anonymous):

yep, can you help me with this one too Each side of a square is increasing at a rate of 9cm/s. At what rate is the area of the square increasing when the area of the square is 36cm2.

OpenStudy (bahrom7893):

sorry im late for college.. all right last one..

OpenStudy (bahrom7893):

A = 36cm^2 => side = 6 cm

OpenStudy (bahrom7893):

A = side^2 dA/dt = 2*s*ds/dt = 2*6cm*9(cm/s) = 108cm^2/s

OpenStudy (anonymous):

when will you be on again. thanks for the help

OpenStudy (bahrom7893):

no clue.. but email me if anything

OpenStudy (anonymous):

oh what is your email?

OpenStudy (bahrom7893):

did u get it?

OpenStudy (anonymous):

it popped up on my screen, how do i go to it?

OpenStudy (bahrom7893):

click notifications

OpenStudy (anonymous):

ok ok i got it

OpenStudy (anonymous):

thanks once again. hope you have a good day

OpenStudy (bahrom7893):

np, u too